# LLM-Enhanced Static Analysis for Precise Identification of Vulnerable OSS Versions

### v1

@@ 3.2.1. AST-level semantic comparison
Input : Post-modification statement’s $AST_{v}$, Pre-modification statements’s $AST_{i}$
Output : Boolean indicating if the commit introduces vulnerabilities
2
3Function ASTSimilarity($AST_{v},AST_{i}^{{}^{\prime}}$):
4       $Sq_{v},Sq_{i}=inOrderTraversal(AST_{v}),inOrderTraversal(AST_{i}^{{}^{\prime}})$
5       $Similarity=EditDistance(Sq_{v},Sq_{i}^{{}^{\prime}})$
6       if $Similarity\geq\theta$ then
7             $return\ True$
8      else
9            $return\ False$
10       end if
11      Function inOrderTraversal($AST$):
12             if $AST\ is\ null$ then
13                   $return\ empty\ list$
14             end if
15            $return\ inOrderTraversal(AST.left)\ +\ [AST.node]\ +\ inOrderTraversal(AST.right)$
16
17
18
Algorithm 1 AST-based Similarity for Code Changes
3) AST similarity. Ultimately, we utilize an in-order traversal algorithm to ascertain whether the commit introduces $S_{v}$ by comparing the AST similarity of $S_{v}$ and $S_{i}^{{}^{\prime}}$, as shown in Algorithm 1. First, we perform an in-order traversal of the ASTs for $S_{v}$ and $S_{i}$, converting the results into sequences $Sq_{v}$ and $Sq_{i}^{{}^{\prime}}$. Then, we calculate the edit distance between the two sequences as their semantic similarity $Similarity_{B}(Sq_{v},Sq_{i}^{{}^{\prime}})$. We set a similarity threshold $\vartheta_{2}$, and if $Similarity_{B}(Sq_{v},Sq_{i}^{{}^{\prime}})\geq\vartheta_{2}$, $S_{v}$ and $S_{i}^{{}^{\prime}}$ are considered to be semanticly similar. If $Similarity_{B}(Sq_{v},Sq_{i}^{{}^{\prime}})<\vartheta_{2}$, it indicates that there is no statement similar to $S_{v}$ in the pre-modification commit.
3.2.2. Code clone determination
We calculate the similarity_score for each code modification of the previous commit with the backtrace. We set a threshold $\vartheta_{3}$, if $similarity\_score<\vartheta_{3}$, we determine that there is no code clone between the pre-modification code of this previous commit and the vulnerable statements, thus confirming that this commit introduced the vulnerability. Otherwise, we proceed with the commit backtrace process.
$\begin{split}similarity\_score=\frac{sim\_a\times weight\_v+sim\_b\times weight\_d}{a\times weight\_v+b\times weight\_d}\end{split}$
where, $a$ and $b$ respectively refer to the counts of all the statements of $S_{v}$ in two different weight types. sim_statement refers to the statements where $Similarity_{A}(S_{v},S_{i}^{{}^{\prime}})\geq\vartheta_{1}$ or $Similarity_{B}(Sq_{v},Sq_{i}^{{}^{\prime}})\geq\vartheta_{2}$ (i.e., semantically or syntactically similar) during each commit comparison. $sim\_a$ and $sim\_b$ respectively refer to the sim_statement counts in two different weight types.
3.3. Vulnerable Version Range Delineation
Since vulnerabilities introduced by vic will be transferred to future commits until they are patched, we identify the versions affected by a vulnerability based on the CVE’s patch commits and the corresponding vic. During the development process of a GitHub project, developers assign tags for different versions. The vulnerable version tags are located between the vic and patch commit.
If commit A is reachable to another commit B, it indicates that A belongs to the parents or the ancestor commits of B (Bao et al., 2022). We first identify the version tags reachable to vic (i.e., $V_{i}$). Then, we identify the version tags reachable to pc (i.e., $V_{p}$). Finally, we obtain the vulnerable version set $V_{v}$, where $V_{v}=V_{i}-V_{p}$. For example of the CVE-2021-20294 (CVE, 2021a), the version tags reachable to pc are [binutils-2_36, binutils-2_42], and the version tags reachable to vic are [binutils-2_35, binutils-2_42], binutils-2_42 is the latest version tags of Binutils project. Therefore, we consider the version tags [binutils-2_35, binutils-2_36) as vulnerable.
4. Experimental Setup
4.1. Supporting Tools
We utilized Joern (Joe, 2014) for program slicing and developed Python scripts to extract control and data dependencies, outputting the sliced statements as dangerous flows. We explored two LLM models, GPT-4 and CodeLlama-13B, for vulnerable code extraction. We use the public API developed by OpenAI to perform the experiment in GPT-4. We set the temperature parameter to 0 to obtain conservative and deterministic answers and the maximum size of tokens to 1024. We deployed the Codellama model on our server with four NVIDIA RTX A5000 GPUs. Similar to GPT-4, the temperature parameter is set to 0, and the maximum size of tokens is set to 1024. Subsequently, we apply weight allocation strategies to extract vulnerable statements using Python scripts. In the code changes detection, we utilize the Clang 10.0.0 tool for function expansion (inline optimization) and AST generation of C/C++ code. Specifically, we use the -fsyntax-only option to perform syntax checking without actual compilation. We then develop Python scripts to achieve AST normalization and similarity comparison. Overall, we constructed our system with 6K LoC in Python.
4.2. Dataset
With confirmation from the author of V-SZZ, it is established that V-SZZ assumes the deleted lines in the patch commit as vulnerable codes and manually labels the first introduction of deleted lines in a patch commit as the vic to construct the dataset. However, as mentioned in Section 2.3, deleted lines cannot fully capture the vulnerability logic. Additionally, it cannot be guaranteed that the identified software version will necessarily exhibit the vulnerability. Therefore, we need to construct a ground-truth dataset with labeled vulnerable version or non-vulnerable version of OSS, which are verified by public PoC.
Vulnerabilities. Our dataset comprises 74 publicly disclosed CVEs associated with public PoC and patches, containing 11 different C/C++ OSS in Table 2. These CVEs span 11 prevalent OSS, as detailed in Table 2. The selected OSS is widely used and has a certain number of versions. Meanwhile, the collected vulnerabilities cover 13 common CWE types.
Patches. We acquired security patches by crawling patch commits from the OSS Git repositories. In total, the patches cover 1,357 modification lines containing 941 insertion lines and 416 deletion lines. Among these, 246 lines are unrelated to vulnerability (e.g., modifications in the ChangeLog file). As show in Table 2, on average, each patch commit has 21.22 modification lines. The number of added lines significantly exceeds that of deleted lines, underscoring the critical importance of incorporating additional code into the analysis process. Among these, 12.5% are large patch commits (the number of modification lines exceeding 50). Notably, 47 patches (63.51%) contain both code insertion and deletion, 21 patches exclusively insertions (28.38%), and 6 patches solely involve deletions (8.11%).
1) The fixed version can be determined through the official release of patch commits. We confirm whether versions following the fixed version are vulnerable by checking if they contain the patch code.
2) For the remaining version, we obtained the public PoC of each vulnerability from the Git issues or referring to OSS sites.
3) To validate the PoC on each corresponding OSS version, we establish the dependent environment for each OSS version and compile the libraries.
4) During the PoC input testing process for each vulnerability, we analyze the triggering conditions and dangerous behaviors of the vulnerability and label the statements that trigger the vulnerability as $S_{trigger}$. For versions where the PoC cannot trigger the vulnerability (as some PoCs are only applicable to certain specific versions), we check whether the version contains $S_{trigger}$ to label the version.
Table 2. Ground-truth Dataset Overview.
IDX Name #Version #CVE #Average Diff Lines Domain
1 Binutils 139 11 14.64 (+12.09/-2.55) Programming tools
2 cJSON 46 1 30.00 (+29.00/-1.00) JSON parser
3 FFmpeg 382 18 21.72 (+12.72/-9.00) Multimedia processing
4 Libarchive 42 9 16.83 (+14.00/-0.86) Streaming processing
5 Libcaca 7 3 36.67 (+25.67/-11.00) Graphics library
6 Liblouis 62 6 4.67 (+3.17/-1.50) Braille translator
7 Libming 14 9 12.22 (+9.78/-2.44) Flash library
8 Libtiff 75 6 16.17 (+13.50/-2.67) Image tools
9 Libxml2 204 3 43.00 (+18.75/-24.25) XML toolkit
10 OpenJPEG 22 5 17.80 (+13.40/-4.40) Image codec
11 Pcre2 20 3 21.67 (+15.00/-6.67) Regular expression engine
Total - 1013 74 21.22 (+15.19/-6.03) -
4.3. Baseline and Metrics
Baseline. For the vulnerable version identification task, we selected the NVD (CPE, 2009), along with SOTAs: SZZ algorithms (AG-SZZ (Kim et al., 2006), B-SZZ (Falleri et al., 2014), V-SZZ (Bao et al., 2022)) and V0Finder (Woo et al., 2021) for accuracy comparison. We gathered the vulnerable version range of each CVE from CPE to evaluate the NVD’s accuracy. The basic idea of SZZ algorithms is to backtrace the commit history to locate the earliest commit that introduces the deletion lines removed by security patches. We use different SZZ algorithms to identify the vic, with the versions between the vic and the pc considered vulnerable. Since these algorithms trace deleted lines individually, instances may arise where multiple vic are identified, and in such cases, we select the earliest one as the detection outcome. V0Finder is a vulnerability introduction detection tool that uses a clone-based technique. We first use V0Finder to generate vulnerable function fingerprints for our ground-truth dataset, and then generate fingerprints for all functions of all OSS versions. Subsequently, we use V0Finder’s clone-based approach to identify vulnerable versions. For Vercation, We selected thresholds $\vartheta_{1}=0.9$, $\vartheta_{2}=0.8$ and $\vartheta_{3}=0.7$ (related experiments are introduced in Section 5.3).
In terms of the performance of vulnerability code extraction, we compared the differences between combining various LLMs (GPT-4, CodeLlama-13B) and using the static analysis tool Joern parser (Joe, 2014) alone.
Metrics. To evaluate the accuracy, we employ the 3 following three metrics, i.e., true positives (TP), false positives (FP), false negatives (FN), precision (#$\frac{\#TP}{\#TP+\#FP}$), recall (#$\frac{\#TP}{\#TP+\#FN}$) and F1-score ($\frac{2*precision*recall}{precision+recall}$), to measure the accuracy of the above methodologies. These metrics are also used in previous studies to evaluate the technique performance (Bao et al., 2022; Rosa et al., 2021). Because of significant differences in the version count for distinct OSS, we compute the precision and recall for each OSS and report the average of precisions and recalls finally.
5. Evaluation
In this section, we answer the following research questions to evaluate the performance of the proposed Vercation.
RQ1. Is Vercation more effective than the state-of-the-art vulnerable version identification methods?
RQ2. How effective is Vercation powered by LLMs in vulnerability comprehension?
RQ3. How accurate is our proposed semantic-level code clone detection method?
RQ4. What are the effects of applying Vercation in the real world?
5.1. RQ1: Effectiveness of Vercation
Table 3 shows the vulnerable version identification results of NVD, SZZ algorithms, V0Finder and Vercation. Note that some CVEs do not provide vulnerable versions in NVD, such as CVE-2021-30499 (CVE, 2021b). Meanwhile, SZZ algorithms cannot work on the patch commits containing solely added lines (i.e., 11 cases), so we exclude these cases. In the Vulnerable Code Extraction phase of Vercation, we utilize the optimal LLM (GPT-4) along with the most effective prompt strategy (Few-shot+CoT), which are comprehensively detailed in Section 5.2. Vercation identified 1612 correct vulnerable versions with 89.2% precision and 95.8% recall, outperforming than the other SOTAs.
5.1.1. Comparison with NVD
NVD reports 224 true positives, 333 false positives and 1398 false negatives with 69.9% precision and 42.4% recall.
1) The reason for false positives in NVD is that it sometimes assumes that all previous versions of disclosed vulnerabilities are vulnerable, without conducting a specific analysis. This situation exists in 12 CVEs (16%) of the ground truth. For instance, security experts found a buffer overflow risk exists in Libtiff 4.4.0, but NVD simply reported the versions of Libtiff$<$4.4.0 were exposed in vulnerability. In reality, the vulnerable version range is Libtiff$<$4.4.0 and $\geq$4.0.0.
2) The reason for low recall in NVD can be summarized into two types: i) NVD only reported the version in which vulnerabilities were found. For example, CVE-2017-14152 (CVE, 2017a) was found in OpenJPEG 2.2.0, then NVD reported the affected version was only 2.2.0, whereas the real vulnerable version range was 2.2.1-2.2.0. The problem exists in 45 CVEs (61%) of the ground truth. ii) Some vulnerabilities are not fixed promptly after disclosure but are addressed after several versions have been released. These vulnerable versions can be overlooked if NVD does not update the version information. For example, CVE-2021-33815 (CVE, 2021c) was discovered in FFmpeg 4.4 but was fixed in version 5.0. Due to the lack of timely tracking of the fix information, NVD overlooked the vulnerable versions 4.4.1-4.4.4.
Table 3. Accuracy of Vulnerable Version Identification.
Methods Reference Correctvul Precison Recall F1 score
AG-SZZ  (Fan et al., 2019) 962 0.405 0.627 0.492
B-SZZ  (Falleri et al., 2014) 901 0.373 0.550 0.445
V-SZZ  (Bao et al., 2022) 1169 0.782 0.889 0.832
NVD  (NVD, 2009) 224 0.699 0.424 0.528
V0Finder  (Woo et al., 2021) 1374 0.834 0.745 0.787
Vercation Submission 1612 0.892 0.958 0.924
5.1.2. Comparison with SZZ algorithms
The basic idea of SZZ algorithms is to backtrace the commit history to locate the earliest commit that introduces the deletion lines removed by security patches. Regarding precision, recall, and F1-score, Vercation improves the best-performing baseline V-SZZ by 11.0%, 6.9% and 9.2%. And the difference between the F1-score of B-SZZ and AG-SZZ is small (0.413 vs. 0.492). We affirmed that there are two main causes of false alarms in this approach:
i) The line mapping algorithm based on edit distance does not consider the semantics of the source code. As shown in Listing LABEL:list:false_reason2, The deleted lines were encapsulated within the added functions, and in fact, they have the same semantics. Line mapping algorithms failed in these cases, unable to identify the true vic.
ii) The SZZ algorithm treats each deleted line as an independent origin for backtracking, resulting in multiple vic being identified. For example, the security patch of CVE-2020-35965 (cve, 2021) from FFmpeg has 5 deleted lines. The true vulnerability logic lies in the lack of checking the size of ymax before executing the memset zero operation, potentially resulting in out-of-bounds writes to memory. Our method utilizes a weighted allocation approach to increase the weight of memset operation, thus backtracing to the correct vic. However, SZZ algorithms separately backtrace different deleted lines, resulting in the identification of 5 different vic, significantly reducing the precision.
5.1.3. Comparison with V0Finder
In the same dataset, V0Finder achieved a precision of 83.4% and a recall of 74.5%. There are two reasons for the inaccuracies of V0Finder: 1) V0Finder generates a hash value for the entire vulnerable function as a vulnerability fingerprint. This fingerprint is coarse because the vulnerable function contains a lot of vulnerability-irrelevant code, introducing a significant amount of noise. 2) V0Finder considers the function containing all deleted patch code as a vulnerable function clone. This approach also overlooked the inserted code of security patch.
Accuracy of Vercation. Though Vercation performed well in the version identification, it still failed to identify inducing commits for a small number of vulnerabilities. One reason is that in some cases, slight modifications in the AST have little impact on the similarity between the AST before and after the modification, but are crucial in introducing vulnerabilities. Another reason is that despite using the weight allocation strategies, there are still some statements related to patch variables that are not relevant to the vulnerability but have been assigned high weights. The above two reasons may incorrectly identify vic, leading to a misidentification of the vulnerable version range.
RQ1: Vercation demonstrates superior performance in vulnerable version identification on our ground-truth dataset, achieving an F1 score of 92.4%. This significantly outperforms existing methods, including NVD reports, SZZ algorithms, and V0Finder. Compared to the best-performing baseline V-SZZ, Vercation shows a 9.2% improvement. These results highlight Vercation’s effectiveness in addressing limitations of current approaches, providing a more accurate tool for identifying vulnerable versions in open-source software.
5.2. RQ2: Ablation Study
We conduct ablation experiments to investigate the effectiveness of being harnessed by LLMs in terms of vulnerability comprehension task and compare the performance with only using the static analysis tool Joern parser.
Table 4. Performance of LLMs in Vulnerability Logic Comprehension.
Tool/Model Prompt Strategy Precision Recall F1 score
Joern - 0.426 0.574 0.489
Joern+GPT-4 Zero-shot 0.708 0.813 0.757
Joern+GPT-4 Few-shot 0.875 0.926 0.899
Joern+GPT-4 Few-shot+CoT 0.892 0.958 0.924
Joern+CodeLlama Zero-shot 0.683 0.780 0.728
Joern+CodeLlama Few-shot 0.863 0.911 0.886
Joern+CodeLlama Few-shot+CoT 0.868 0.938 0.901
5.2.1. Compared with Only Using Joern parser
As depicted in Table  4, the effectiveness of Joern in extracting vulnerability logic is significantly inferior to hardnessed by the power of LLM models, with an F1 score 45.4% less than Joern+GPT-4 (Few-shot+CoT). This discrepancy primarily stems from Joern’s reliance on fixed dependency extraction patterns, leading to the retrieval of numerous statements irrelevant to vulnerabilities. Meanwhile, we found that Joern cannot accurately address pointer structure dataflows, thus missing some statements that have data or control dependencies. To elaborate, on average, Joern extracts 3.12 and 2.71 times more vulnerable statements per CVE compared to combined GPT-4 and Codellama, respectively.
5.2.2. Performance of different LLM models
Table  4 illustrates the performance of combining the dangerous flows extracted by the Joern parser with GPT-4 and CodeLlama-13B models on our dataset, with the ”Prompt” column delineating different prompting strategies. Our observation reveals that GPT-4 prompted with Few-shot and CoT surpasses other models. Specifically, GPT-4 (Few-shot+CoT) exhibits 2.3% higher F1 score compared to the CodeLlama model. Moreover, the performance trends with GPT-4 remain consistently upward across different prompting strategies. When prompted with the Few-shot prompt, GPT-4 achieves a 14.2% higher F1 score than the Zero-shot prompt. This is further improved by 2.5% F1 with the addition of CoT prompting. We observed utilization of the Zero-shot prompt often results in an increased generation of false positive vulnerable statements, including the statements predefining patch variables (e.g., Line 9 in Listing 1). Conversely, in examples prompted by Few-shot, we prioritize understanding the triggering mechanisms of vulnerabilities, thereby facilitating LLM in learning a more accurate method of generating vulnerability logic.
RQ2: Harnessed by the power of LLMs, Vercation can generate the vulnerability logic and extract vulnerable codes from dangerous flows more precisely than traditional static analysis tools. The prompt strategy combining Few-shot and CoT techniques performs the best.
5.3. RQ3: Accuracy of Code Clone Detection
Here we evaluate the accuracy of the semantic-level similarity comparison technique utilized in P2.
5.3.1. Threshold Sensitivity
We used $\vartheta_{1}$ and $\vartheta_{2}$ to represent the threshold for syntactic similarity and semantic similarity respectively. $\vartheta_{3}$ represents the similarity_score used for the backtrace termination condition. To measure the sensitivity of the thresholds, we incrementally increased $\vartheta_{1}$, $\vartheta_{2}$ and $\vartheta_{3}$ by 0.1 from 0 to 1, and evaluated the identification results of each CVE in ground-truth. In addition, we evaluated each threshold with the metric F1-score, as we focused on both precision and recall. As shown in the results of Figure 4, the highest accuracy is achieved when $\vartheta_{1}$ is 0.9, $\vartheta_{2}$ is 0.8 and $\vartheta_{3}$ is 0.7. Specifically, when the thresholds are greater than the optimal values, the precision is higher while the recall decreases. In contrast, the precision slightly decreases.
Figure 4. Threshold Sensitivity of Vercation.
5.3.2. Accuracy of Semantic-level Code Clone Detection
Current methods primarily focus on function and file-level analysis[6,7], using coarse-grained representations like identifier sequences and control-flow graphs. While Vercation operates at sentence level, providing more nuanced analysis, differs significantly from existing code clone detection techniques. These conventional methods are not well-suited for sentence-level scenarios. Therefore, we compare some typical sentence-level code similarity detection methods with the method in Vercation’s P2. Table 5 summarizes the accuracy of different code clone detection methods.
1) Edit distance comparison calculates the minimum number of edit operations (insertions, deletions, substitutions) required to transform one code sentence into another. A smaller edit distance indicates a higher similarity between two code sentences.
2) Hash values comparison generates hash values for each code sentence and then compares the similarity of these hash values. Identical or similar hash values suggest that the code sentences may be similar.
3) CodeBERT embedding comparison uses the CodeBERT to generate embedding representations for each code sentence, then compute the cosine similarity between these embeddings. A cosine similarity closer to 1 indicates a higher semantic similarity between the code sentences.
[…]

### v2

@@ V-B1 Program Slicing & LLM Ablation Study
    The average count of extracted vulnerable statements for each CVE.
Static Analysis Only. As depicted in Table IV, the effectiveness of Joern in extracting vulnerability logic is significantly inferior to that harnessed by the power of LLM models, with an F1 score 46.9% less than Joern + DeepSeek (Few-shot + CoT). This discrepancy primarily stems from Joern’s reliance on fixed dependency extraction patterns, leading to the retrieval of numerous statements irrelevant to vulnerabilities. As shown in Figure 5(a), using only the static analysis tool Joern parser for program slicing extracts statements on Lines 3, 7-8, 11-12. While these statements are related to control flow, they are not directly relevant to the vulnerability. This illustrates how static analysis alone can include extraneous information that is not crucial to understanding the specific vulnerability logic. Meanwhile, we found that Joern cannot accurately address pointer structure dataflows, thus missing some statements that have data or control dependencies. To elaborate, on average, Joern extracts 3.64 times more vulnerable statements per CVE compared to combined DeepSeek.
LLM Only. The approach of LLM Only (DeepSeek) involves inputting the entire function into the model for analysis. As illustrated in Table 5(b), DeepSeek alone (with Few-shot + CoT strategies) achieves an F1 score of 0.739, which is 18.7% lower than the Joern + DeepSeek combination (0.926). This performance gap primarily stems from LLM’s tendency to include a broader range of contextual information when given an entire function as input, this over-inclusion of context leads to lower precision because some of which are tangential to the core vulnerability logic. As illustrated in Figure 5(b), using only LLM improves upon static analysis by reducing the extraction of irrelevant control flow statements. However, when given the entire function as input, the LLM struggles to precisely differentiate between core vulnerability-related code and contextual information. This results in the inclusion of some noise, such as variable definitions (Line 1) and unrelated control flow statements (Lines 4, 7-8).
As shown in Figure 5(c), the combination of Joern’s structured analysis with DeepSeek’s code understanding allows for a more focused and accurate extraction of vulnerability-related statements. Joern provides an initial set of potentially relevant statements based on program structure, which DeepSeek then refines using its contextual understanding, resulting in a more balanced and effective approach to vulnerability logic comprehension.
V-B2 LLM Performance, Consistency, and Generalization
We further evaluated the LLM component by comparing different models and analyzing their stability and ability to generalize.
TABLE V: Vercation with Different Strategies.
Model Strategy Precision Recall F1 score
GPT-4 Zero-shot 0.708 0.823 0.761
Few-shot 0.892 0.907 0.899
Few-shot + CoT 0.893 0.946 0.925
CodeLlama Zero-shot 0.671 0.789 0.725
Few-shot 0.851 0.895 0.873
Few-shot + CoT 0.842 0.912 0.876
DeepSeek-V3 Zero-shot 0.769 0.827 0.797
Few-shot 0.882 0.921 0.901
Few-shot + CoT 0.918 0.945 0.931
Comparative Performance. Table V illustrates the performance of combining the dangerous flows extracted by the Joern parser with LLMs on our dataset, with the “Strategy” column delineating different prompting strategies. Our observation reveals that DeepSeek-V3 prompted with Few-shot and CoT surpasses other models. Specifically, DeepSeek-V3 (Few-shot + CoT) exhibits a 0.64% and 6.28% improvement in F1-score compared to the GPT-4 and CodeLlama, respectively. Moreover, the performance trends with DeepSeek remain consistently upward across different prompting strategies. When prompted with the Few-shot prompt, DeepSeek achieves a 13.15% improvement in F1-score over the Zero-shot prompt. This is further improved by 3.33% F1 with the addition of CoT prompting. We observed that the utilization of the Zero-shot prompt often results in an increased generation of false positives (vulnerable statements), including the statements predefining patch variables (e.g., Line 6 in Listing LABEL:list:CVE-2017-14169). Conversely, in examples prompted by Few-shot, we prioritize understanding the triggering mechanisms of vulnerabilities, thereby facilitating LLM in learning a more accurate method of generating vulnerability logic.
TABLE VI: LLM Consistency for Vulnerable Statement Extraction.
Model Agreement Rate Jaccard Similarity
GPT-4 0.454 0.439
CodeLlama 0.517 0.566
DeepSeek-V3 0.586 0.629
LLM Consistency Analysis. To evaluate the stability of our LLM-based vulnerable statement extraction, we measured two consistency metrics across 10 independent runs with temperature = 1.0. The agreement rate represents the percentage of statements that appeared in all 10 runs relative to the total unique statements extracted across all runs. As shown in Table VI DeepSeek-V3 achieved the highest agreement rate 58.6%, followed by CodeLlama (51.7%). For cases with variation, we calculated average Jaccard similarity by computing pairwise similarities between all run combinations and taking the mean. DeepSeek-V3 achieved the highest Jaccard similarity of 0.629 shown in Table VI, indicating that DeepSeek-V3 maintains the most consistent performance even when exact matches were not achieved. While the consistency levels are moderate across all models, this reflects the inherent complexity of vulnerability logic analysis where slight variations in reasoning can lead to different but potentially valid interpretations. The superior consistency of DeepSeek-V3, combined with its highest F1-score performance, makes it the most reliable choice for practical deployment.
Generalization Capability. A critical concern regarding the LLM-based approach is whether the models’ knowledge cutoff affects performance when analyzing vulnerabilities published after the training data cutoff. To address this concern and demonstrate that Vercation leverages LLMs’ general code understanding capabilities rather than specific vulnerability knowledge, we conducted an ablation study comparing each model’s performance on CVEs published before and after their respective knowledge cutoff dates. The three LLMs in our evaluation have different knowledge cutoff dates: CodeLlama’s training data extends to September 2022 (based on Llama 2’s cutoff date [62]), GPT-4’s knowledge cutoff is December 2023 [63], and DeepSeek-V3’s knowledge cutoff is July 2024 (confirmed through direct inquiry with DeepSeek). We partitioned our dataset according to each model’s cutoff date and evaluated their performance on both pre-cutoff and post-cutoff CVE subsets as shown in Figure 6.
As shown in Table VII, CodeLlama shows a slight performance decrease of 1.25% (from 0.881 to 0.870) when analyzing post-cutoff CVEs. In contrast, both GPT-4 and DeepSeek-V3 demonstrate slight performance improvements on post-cutoff CVEs, with GPT-4 improving by 0.54% and DeepSeek-V3 by 1.73%. The performance variations across all models are minimal (<< 2%), indicating remarkable consistency regardless of knowledge cutoff boundaries. These results validate our core hypothesis that Vercation’s effectiveness stems from leveraging LLMs’ fundamental code understanding and reasoning capabilities rather than memorized vulnerability-specific knowledge.
Figure 6: Temporal Distribution of CVEs with Knowledge Cutoff Boundary.
TABLE VII: Performance Comparison Across Knowledge Cutoff Periods.
Model Cutoff Date Period F1-score
CodeLlama 2022.09 Pre-cutoff 0.881
Post-cutoff 0.870 ($\downarrow$ 1.25%)
GPT-4 2023.10 Pre-cutoff 0.923
Post-cutoff 0.928 ($\uparrow$ 0.54%)
DeepSeek-V3 2024.07 Pre-cutoff 0.927
Post-cutoff 0.943 ($\uparrow$ 1.73%)
Figure 7: Impact of Statement Weight Ratios on F1-Score
V-B3 Effectiveness of the Statement Weighting Strategy
To validate the effectiveness of our statement weighting strategy, we conducted an ablation study comparing our weighted approach with different weight ratios across our entire dataset. We systematically evaluated weight ratios from 1:1 to 5:1, where the first number represents the weight for sensitive function statements ($W_{sensitive}$) and the second number represents the weight for regular statements ($W_{base}$). These ratios were tested using the DeepSeek-V3 model, with all other parameters held constant.
Figure 7 illustrates the impact of different weight ratios on F1-score. The results demonstrate that equal weighting (1:1 ratio) achieves an F1-score of 0.896, while our proposed 2:1 ratio ($W_{sensitive}$ = 2.0, $W_{base}$ = 1.0) achieves the optimal performance of 0.931, representing a 3.91% improvement. Performance gradually decreases as the weight ratio increases beyond 2:1. Higher ratios (3:1 and above) lead to over-emphasis on sensitive functions. This causes the similarity score to be dominated by a small number of statements, reducing the method’s ability to distinguish between true vulnerability patterns and coincidental sensitive function usage.
RQ2: Vercation’s effectiveness stems from the synergistic combination of static analysis and LLMs. DeepSeek-V3 with a Few-shot+CoT prompt proves to be the most accurate, consistent, and generalizable model. Furthermore, our 1:2 statement weighting strategy is shown to be optimal for balancing the influence of sensitive functions.
V-C RQ3: Component Robustness
Here we evaluate the accuracy of the similarity comparison technique utilized in Vercation’s phase 2.
V-C1 Accuracy of Code Clone Detection
We compare the code clone detection technique in Vercation’s phase 2 with typical statement-level code similarity detection methods: Edit distance [64], Hash value [65], and CodeBERT embedding comparison [66]. Current SOTA code clone detection methods primarily focus on function-level detection [18], whereas our method operates at the statement level. To perform a comparison, during the process of backtracing the previous commits of vulnerable statements, we apply code clone detection to the entire function before and after the modification. If the commit includes changes to other non-vulnerable statements within the function, we normalize these statements to be identical. Table VIII summarizes the accuracy of different code clone detection methods.
Edit distance [64] calculates the minimum number of edit operations required to transform one code statement into another. A smaller edit distance indicates a higher similarity. Hash values comparison [65] generates hash for each code statement through MD5 algorithm and then compares the hash similarity. CodeBERT embedding comparison [66] uses CodeBERT to generate embedding representations for each code statement, then computes the cosine similarity between these embeddings. FCDetection [67] generates AST and CFG representations as features. Word2vec and Graph2vec are used to embed these features. Then the fused feature vectors are input into a deep neural network model for classification of code clones.
TABLE VIII: Accuracy of Code Clone Detection Methods.
Methods Level Precision Recall
Edit distance statement 0.85 0.84
Hash values statement 0.82 0.82
CodeBERT statement 0.84 0.88
FCDetector function 0.89 0.93
Vercation statement 0.90 0.95
For Edit distance, Hash values, and CodeBERT embedding comparisons, we conducted independent threshold optimization. Similar to our threshold sensitivity analysis for Vercation, we evaluated each method across different threshold values from 0.1 to 1.0 with a step of 0.1, and selected the threshold that achieved the best F1 for that specific method. As FCDetector is a specialized code clone detection tool, we used its original pre-trained model and thresholds as specified in their work. The results presented in Table VIII represent each method’s optimal performance with their respective optimized thresholds. The results show that the AST-based code clone detection achieves the best performance. Because the edit distance and hash value methods terminate the backtrack in cases of low textual similarity without considering code behavior, thus leads to a large number of false negatives. The results show that our method is robust against low syntactic similarity. CodeBERT embedding comparison, while effective for semantic similarity, cannot capture fine-grained structural information. AST representations preserve the exact syntactic structure of code, allowing for more precise comparisons of code organization and logic flow. This structural analysis can identify clones that CodeBERT might miss, especially in cases where similar functionality is implemented with different coding patterns or variable names, which can be addressed by AST normalization. Although FCDetector also extracts AST as a feature and normalizes variable names and constant values, it cannot identify function outline cases (as shown in Listing LABEL:semantic). Moreover, the CFG features that FCDetector focuses on are not particularly effective when there are only a few vulnerable statements.
V-C2 Analysis of Refactoring Commits
As motivated in Example 2 of Section II-C, refactoring commits can significantly impact vulnerable version identification by introducing structural differences while preserving code behavior. To validate this motivation and evaluate the effectiveness of our code clone detection approach, we analyzed the commits encountered during our evaluation. In the backtrack of our dataset, we conducted a total of 276 comparisons between pre-commit and post-commit. On average, it takes 3.73 comparisons to identify the vic for each vfc. During the commit backtrack, cases with significant edit-distance differences ($\theta_{1}<0.9$) but minor AST differences ($\theta_{2}\geq 0.8$) were classified as refactoring commits. Out of these comparisons, 89 commits (32.2%) were identified as refactoring commits, indicating that refactoring commits are a significant factor in backtracing vic. Furthermore, our analysis reveals that cases without any intermediate commit between the vic and the vfc are comparatively uncommon, constituting only 11% of instances in our dataset.
V-C3 Threshold Sensitivity
We used $\vartheta_{1}$ and $\vartheta_{2}$ to represent the threshold for edit-distance similarity and AST-based similarity, respectively. $\vartheta_{3}$ represents the similarity_score used for the backtrack termination condition. To measure the sensitivity of the thresholds, we incrementally increased $\vartheta_{1}$, $\vartheta_{2}$ and $\vartheta_{3}$ by 0.1 from 0 to 1, and evaluated the identification F1-score of the CVEs in ground-truth. Specifically, we first initialized $\vartheta_{2}$ and $\vartheta_{3}$ to 0.8, as high similarity thresholds generally provide better precision in code clone detection. Then we varied $\vartheta_{1}$ from 0.1 to 1.0 with a step of 0.1 and found the value of $\vartheta_{1}$ that achieved the highest F1-score (0.9). With $\vartheta_{1}$ fixed at 0.9 and kept $\vartheta_{3}$ at 0.8, we varied $\vartheta_{2}$ from 0.1 to 1.0 and found the optimal value (0.8). Finally, with $\vartheta_{1}$ = 0.9 and $\vartheta_{2}$ = 0.8, we varied $\vartheta_{3}$ from 0.1 to 1.0 and determined the optimal value of $\vartheta_{3}$ (0.7). When the thresholds are greater than the optimal values, the precision is higher while the recall decreases. In contrast, the precision slightly decreases.
Figure 8: Threshold Sensitivity of Vercation.
RQ3: Refactoring commits occupy a significant proportion of the code changes history. The AST-based similarity comparison in Vercation performs better in the vulnerability-introducing commit identification task for calculating code similarity, enhancing the overall effectiveness of Vercation in vulnerable versions detection.
V-D RQ4: Real-World Application
We created a new vulnerability dataset consisting of 342 CVEs from the NVD. By developing a simple crawler, we obtained the corresponding CPE and CVE patches. We applied Vercation to the vulnerability dataset and found the vulnerable versions of 202 CVEs are incorrect. Among them, the CPE for 134 CVEs is incomplete, with 108 CVEs only reporting a single vulnerable version, significantly increasing the risk of vulnerability propagation. Furthermore, the CPE for 68 CVEs is affected by the overinclusion problem, with 56 CVEs considered all versions before the OSS version mentioned in the vulnerability report as susceptible to attack, without conducting a detailed analysis. We have submitted reports to the NVD detailing the CVEs with incorrect affected versions information we identified. The NVD team’s response said they plan to address these inaccuracies as part of their ongoing efforts to enhance the quality of their vulnerability information. More importantly, we have discovered 4 CVEs that do not provide information about the vulnerable version.
[…]